OpenStudy (anonymous):

Choose the fraction that goes in the blank. 5/12<___<3/4 A. 5/6 B. 1/4 C. 2/3 D. 7/8

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }*\frac{ 3 }{ 3 }=\frac{ 9 }{ 12 }\]

OpenStudy (texaschic101):

5/12 < ___ < 3/4 --- convert to decimals 0.416 < ___ < 0.75 so we know that the missing number has to be between 0.416 and 0.75. So we will look at our answer choices and convert each one to a decimal 5/6 = 0.833 1/4 = 0.25 2/3 = 0.666 7/8 = 0.875 so what do you think ?